OpenStudy (anonymous):

Find the point (x,y) on the graph of y=rootx that is closest to the point (4,0). Please help!

jimthompson5910 (jim_thompson5910):

The equation is \[\Large y = \sqrt{x}\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Any point (x,y) on this graph is going to be the point \(\Large \left(x, \sqrt{x}\right)\). Agreed? or no?

OpenStudy (anonymous):

okay, got it

jimthompson5910 (jim_thompson5910):

so here is how I would approach it

jimthompson5910 (jim_thompson5910):

First point (fixed) \[\Large P = (x_1,y_1) = (4,0)\] \[\Large x_1 = 4\] \[\Large y_1 = 0\] Second point (moving) \[\Large Q = (x_1,y_1) = \left(x, \sqrt{x}\right)\] \[\Large x_1 = x\] \[\Large y_1 = \sqrt{x}\] where x is nonnegative Distance from point P to point Q (using the distance formula) \[\Large d(P,Q) = \sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\] \[\Large d(P,Q) = \sqrt{(4-x)^2+(0-\sqrt{x})^2}\] \[\Large d(P,Q) = \sqrt{(4-x)^2+(-\sqrt{x})^2}\] \[\Large d(P,Q) = \sqrt{16-8x+x^2+x}\] \[\Large d(P,Q) = \sqrt{x^2-7x+16}\] The function \[\Large d(x) = \sqrt{x^2-7x+16}\] represents the distance from point P (4,0) to point Q (any point on \(\Large y = \sqrt{x}\)) The goal is to minimize d(x)

jimthompson5910 (jim_thompson5910):

Let me know if you're able to minimize d(x)

ganeshie8 (ganeshie8):

you could also try minimizing \(d^2\), the extrema of \(|d|\) and \(|d|^2\) occur at the same \(x\) value

OpenStudy (anonymous):

okay, would it be x^2-7/2rootx^2-7x+16?

jimthompson5910 (jim_thompson5910):

\[\Large d(x) = \sqrt{x^2-7x+16}\] \[\Large d \ ' (x) = \frac{2x-7}{2\sqrt{x^2-7x+16}}\]

jimthompson5910 (jim_thompson5910):

you should have 2x-7 up top (not x^2-7)

jimthompson5910 (jim_thompson5910):

Now solve \[\Large d \ ' (x) = 0\]

OpenStudy (anonymous):

would my point be (7/2,rott7/2)?

jimthompson5910 (jim_thompson5910):

x = 7/2 is correct idk how you got \(\Large \sqrt{\frac{7}{2}}\) for the distance

jimthompson5910 (jim_thompson5910):

oh nvm I'm not thinking

jimthompson5910 (jim_thompson5910):

yes, the point \[\Large \left(\frac{7}{2}, \sqrt{\frac{7}{2}}\right)\] is the right final answer

OpenStudy (anonymous):

okay! thank you!